# The Analogue of Aldous’ spectral gap conjecture for the generalized exclusion process

### v2

@@ 3.2. Generalized exclusion process
For any $l\in\{1,\dots n\}$, let $H_{l}$ be a subgroup of $\mathfrak{S}_{n}$ corresponding to the permutation group of $\{1,\dots,l\}$, and let $H_{l,n}$ be a subgroup of $\mathfrak{S}_{n}$ with corresponding to the permutation group of $\{l+1,\dots,n\}$. Then, by Lemma 3.3, $H_{l}$ and $H_{l,n}$ naturally acts on $\mathrm{IP}(X)$. We can regard $S_{\mathrm{IP}}/H_{l}$ as the set of assignments of $n-l$ labeled particles to the vertices of $X$ in such a way that each vertex is occupied by exactly one particle. We can regard $S_{\mathrm{IP}}/H_{l}H_{l,n}$ as the set of assignments of $l$ indistinguishable particles to the vertices of $X$ in such a way that each vertex is occupied by exactly one particle. Therefore, the complete weighted directed graph $\mathrm{IP}(X)/H_{l}H_{l,n}$ is isomorphic to the symmetric exclusion process on $X$. We set $\mathrm{SEP}(X)=(S_{\mathrm{SEP}},W_{\mathrm{SEP}}):=\mathrm{IP}(X)/H_{l}H_{l,n}$. If $l=1$, we will especially write $\mathrm{SEP}(X)$ as $\mathrm{RW}(X)$. We naturally identify $S_{\mathrm{SEP}}$ with the set
@@
Let $N=kl$, and let $l\in\{1,\dots,N-1\}$. Set $\mathfrak{H}_{l}:=H_{l}H_{l,N}$. We let $\mathrm{SEP}(\widetilde{X})=(\widetilde{S}_{\mathrm{SEP}},\widetilde{W}_{\mathrm{SEP}}):=\mathrm{IP}(\tilde{X})/\mathfrak{H}_{l}$. For any $v\in V$, put $H_{v}\cong\mathfrak{S}_{k}$. We next define an action of $\prod_{v\in V}H_{v}$ on $\mathrm{SEP}(\widetilde{X})$. We define the action of $H_{v}$ on $\widetilde{V}$ by
@@
$\varphi_{GEP}:\widetilde{S}_{\mathrm{SEP}}/\mathbb{H}\rightarrow S_{\mathrm{GEP}}$
by $\varphi_{GEP}(\eta)(u)=\sum_{\widetilde{u}\in\varphi_{V}^{-1}(u)}\eta(\widetilde{u})$.
@@
The map $\varphi_{GEP}$ is a morphism of complete weighted directed graphs. In other words, the generalized exclusion process $\mathrm{GEP}(X)$ is a quotient of the interchange process $\mathrm{IP}(\widetilde{X})$.
@@
$\sum_{\eta_{2}\in\varphi_{GEP}^{-1}(\pi_{2})}\widetilde{W}_{\mathrm{SEP}}(\eta_{1},\eta_{2})=\pi_{1}(u)(k-\pi_{1}(v))r(u,v)=W_{\mathrm{GEP}}(\pi_{1},\pi_{2}).$
@@ 4. Main theorem
$\lambda_{1}^{\mathrm{IP}(X)}=\lambda_{1}^{\mathrm{SEP}(X)}.$
@@
Therefore, we have $\lambda^{\mathrm{RW}(X_{k})}=k\lambda^{\mathrm{RW}(X_{k})}$. ∎
@@
$\lambda_{1}^{\mathrm{GEP}(X)}=k\lambda_{1}^{\mathrm{RW}(X)}.$
@@
For any $i\in\{1,\dots,N\}$, we $H^{i}$ let denote the permutation group of $\{1,\dots,i-1,i+1,\dots,N\}$. For any $h\in H=H_{l}H_{l,N}$, there exists $i=l+1,\dots,N$ such that $hH_{N-1}h^{-1}=H^{i}$. Therefore, we have $\cap_{i=1,\dots,N}H^{i}=H_{l}\subset H_{l}H_{l,N}$. Moreover, the intersection $\cap_{i=1,\dots,N-1}H^{i}$ is the permutation group of $\{1,\dots,l,N\}$. Therefore, the subgroup of $G$ generated by $H_{N-1}$ and $\cap_{i=1,\dots,N-1}H^{i}$ is equivalent to $G$. Hence, if we put $K=H_{N-1}$, $H=H_{l}H_{l,N}$, then $H$ and $K$ satisfies the condition in Theorem 2.16. Therefore, by Theorem 2.16 and Proposition 3.6, we have
$\lambda_{1}^{\mathrm{IP}(\widetilde{X})}\leq\lambda_{1}^{\mathrm{SEP}(\widetilde{X})}\leq\lambda_{1}^{\mathrm{GEP}(X)}\leq\lambda_{1}^{\mathrm{RW}(X_{k})}=k\lambda_{1}^{\mathrm{RW}(X)}.$
On the other hand, by Theorem 4.1 and Theorem 4.2, we have
$\lambda_{1}^{\mathrm{IP}(\widetilde{X})}=\lambda_{1}^{\mathrm{RW}(\widetilde{X})}=\lambda_{1}^{\mathrm{RW}(X_{k})}.$
Therefore, we have
$\lambda_{1}^{\mathrm{GEP}(X)}=\lambda_{1}^{\mathrm{RW}(X_{k})}.$

### v3

@@ 3.2. Generalized exclusion process
For any $l\in\{1,\dots n\}$, let $H_{l}$ be the subgroup of $\mathfrak{S}_{n}$ corresponding to the permutation group of $\{1,\dots,l\}$, and let $H_{l,n}$ be the subgroup of $\mathfrak{S}_{n}$ corresponding to the permutation group of $\{l+1,\dots,n\}$. Then, by Lemma 3.3, $H_{l}$ and $H_{l,n}$ naturally act on $\mathrm{IP}(X)$. We can regard $S_{\mathrm{IP}}/H_{l}$ as the set of assignments of $n-l$ labeled particles to the vertices of $X$ in such a way that each vertex is occupied by exactly one particle. We can regard $S_{\mathrm{IP}}/H_{l}H_{l,n}$ as the set of assignments of $l$ indistinguishable particles to the vertices of $X$ in such a way that each vertex is occupied by exactly one particle. Therefore, the complete weighted directed graph $\mathrm{IP}(X)/H_{l}H_{l,n}$ is isomorphic to the symmetric exclusion process on $X$. We set $\mathrm{SEP}_{l}(X)=(S_{\mathrm{SEP}},W_{\mathrm{SEP}}):=\mathrm{IP}(X)/H_{l}H_{l,n}$. If $l=1$, we will especially write $\mathrm{SEP}_{l}(X)$ as $\mathrm{RW}(X)$. We naturally identify $S_{\mathrm{SEP},l}$ with the set
@@
Let $N=kl$, and let $l\in\{1,\dots,N-1\}$. Set $\mathfrak{H}_{l}:=H_{l}H_{l,N}$. We let $\mathrm{SEP}_{l}(\widetilde{X})=(\widetilde{S}_{\mathrm{SEP}},\widetilde{W}_{\mathrm{SEP}}):=\mathrm{IP}(\tilde{X})/\mathfrak{H}_{l}$. For any $v\in V$, put $H_{v}\cong\mathfrak{S}_{k}$. We next define an action of $\prod_{v\in V}H_{v}$ on $\mathrm{SEP}_{l}(\widetilde{X})$. We define the action of $H_{v}$ on $\widetilde{V}$ by
@@
$\varphi_{\mathrm{GEP}_{k,l}(X)}:\widetilde{S}_{\mathrm{SEP}}/\mathbb{H}\rightarrow S_{\mathrm{GEP}}$
by $\varphi_{\mathrm{GEP}_{k,l}(X)}(\eta)(u)=\sum_{\widetilde{u}\in\varphi_{V}^{-1}(u)}\eta(\widetilde{u})$. The image of this morphism is shown in Figure 5.
Figure 5. The lower part represents the particle jumps in GEP with $k=2$ and $l=2$, while the upper part illustrates an example of corresponding state transitions in SEP on the expanded graph.
@@
The map $\varphi_{\mathrm{GEP}_{k,l}(X)}$ is a morphism of complete weighted directed graphs. In other words, the generalized exclusion process $\mathrm{GEP}_{k,l}(X)$ is a quotient of the interchange process $\mathrm{IP}(\widetilde{X})$.
@@
$\sum_{\eta_{2}\in\varphi_{\mathrm{GEP}_{k,l}(X)}^{-1}(\pi_{2})}\widetilde{W}_{\mathrm{SEP}}(\eta_{1},\eta_{2})=\pi_{1}(u)(k-\pi_{1}(v))r(u,v)=W_{\mathrm{GEP}}(\pi_{1},\pi_{2}).$
@@ 4. Main theorem
$\lambda_{1}^{\mathrm{IP}(X)}=\lambda_{1}^{\mathrm{SEP}_{l}(X)}.$
@@
Therefore, we have $\lambda_{1}^{\mathrm{RW}(X_{k})}=k\lambda_{1}^{\mathrm{RW}(X_{k})}$. ∎
Corollary 4.3.
We have the equality
$\lambda_{1}^{\mathrm{SEP}_{l}(\widetilde{X})}=\lambda_{1}^{\mathrm{RW}(\widetilde{X})}=k\lambda_{1}^{\mathrm{RW}(X)}.$
Proof.
The first equality follows from Theorem 4.1, and the second equality follows from Theorem 4.2. ∎
@@
$\lambda_{1}^{\mathrm{GEP}_{k,l}(X)}=\lambda_{1}^{\mathrm{RW}(X_{k})}.$
@@
For any $i\in\{1,\dots,N\}$, we $H^{i}$ let denote the permutation group of $\{1,\dots,i-1,i+1,\dots,N\}$. For any $h\in H=H_{l}H_{l,N}$, there exists $i=l+1,\dots,N$ such that $hH_{N-1}h^{-1}=H^{i}$. Therefore, we have $\cap_{i=1,\dots,N}H^{i}=H_{l}\subset H_{l}H_{l,N}$. Moreover, the intersection $\cap_{i=1,\dots,N-1}H^{i}$ is the permutation group of $\{1,\dots,l,N\}$.
By Corollary 4.3, it suffices to show that
$\lambda_{1}^{\mathrm{SEP}_{l}(\widetilde{X})}=\lambda_{1}^{\mathrm{GEP}_{k,l}(X)}.$
Since there is a canonical morphism $\mathrm{SEP}_{l}(\widetilde{X})\rightarrow\mathrm{GEP}_{k,l}(X)$ and Lemma 2.17, it is enough to show that
$\lambda_{1}^{\mathrm{GEP}_{k,l}(X)}\leq\lambda_{1}^{\mathrm{SEP}_{l}(\widetilde{X})}.$
By Corollary 4.3, there is an eigenvector $f$ of $-\mathcal{L}^{\mathrm{RW}(X_{k})}$ with the eigenvalue $\lambda_{1}^{\mathrm{SEP}_{l}(\widetilde{X})}$. We regard $f$ as the function $V\cong S_{\mathrm{RW}(X_{k})}\rightarrow\mathbb{R}$. Let $\tilde{f}:\widetilde{V}\rightarrow\mathbb{R}$ be a map induced by pulling back $f$ along the canonical surjection $\tilde{V}\rightarrow V$. Then, by the proof of [P10]*Proposition 3.1, $\tilde{f}$ is an eigenvector of $-\mathcal{L}^{\mathrm{RW}(\widetilde{X})}$ with the eigenvalue $\lambda_{1}^{\mathrm{RW}(\widetilde{X})}$. We let define the function $F:S^{\mathrm{SEP}_{l}(\widetilde{X})}\rightarrow\mathbb{R}$ by
$F(\eta):=\sum_{s\in S_{\eta}}\tilde{f}(s)$
where $S_{\eta}:=\eta^{-1}(1)$ for any $\eta\in S^{\mathrm{SEP}_{l}(\widetilde{X})}$. Then, by [CLR10]*§4.1, $F$ is an eigenvector with the eigenvalue $\lambda_{1}^{\mathrm{SEP}_{l}(\widetilde{X})}$. Moreover, if we define $\overline{F}:S^{\mathrm{GEP}_{k,l}(X)}\rightarrow\mathbb{R}$ by
$\overline{F}(\pi):=F(\eta_{\pi}),$
where $\eta_{\pi}\in S^{\mathrm{SEP}_{l}(\widetilde{X})}$ is an arbitrary lift of $\pi\in S^{\mathrm{GEP}_{k,l}(X)}$ along the canonical surjection $S^{\mathrm{SEP}_{l}(\widetilde{X})}\rightarrow S^{\mathrm{GEP}_{k,l}(X)}$. Since $\tilde{f}$ is constant on each fiber $\widetilde{V}\rightarrow V$, this does not depend on the choice of lifts of $\pi$. Moreover, this is an eigenvector of $-\mathcal{L}^{\mathrm{GEP}_{k,l}(X)}$ with the eigenvalue $\lambda_{1}^{\mathrm{SEP}_{l}(\widetilde{X})}$. Indeed, we have
$\begin{split}-\mathcal{L}^{\mathrm{GEP}_{k,l}(X)}\overline{F}(\pi)&=-\mathcal{L}^{\mathrm{GEP}_{k,l}(X)}(F\circ\varphi_{\mathrm{GEP}_{k,l}(X)})(\eta_{\pi})\\ &=-\mathcal{L}^{\mathrm{SEP}_{l}(\widetilde{X})}F(\eta_{\pi})\\ &=\lambda_{1}^{\mathrm{SEP}_{l}(\widetilde{X})}F(\eta_{\pi})\\ &=\lambda_{1}^{\mathrm{SEP}_{l}(\widetilde{X})}\overline{F}(\pi).\end{split}$
Therefore, we conclude that
$\lambda_{1}^{\mathrm{GEP}_{k,l}(X)}\leq\lambda_{1}^{\mathrm{SEP}_{l}(\widetilde{X})}.$